OpenStudy (anonymous):

How much money will there be in an account at the end of 10 years if $4000 is deposited at 3% interest compounded semi-annually? (Assume no withdrawals are made) The amount after 10 years will be $___

OpenStudy (anonymous):

FV = PV(1+ i/2)^n(2) FV = 4000(1+ 0.03/2)^20 FV = 5387.42

OpenStudy (anonymous):

Thanks again. :)

OpenStudy (anonymous):

np